Grade 12Physical Chemistry

the latent heat of vapourisation of a liquid at 500 K and 1 atm pressure is 10 kcal/mol. what will be the change in internal energy of 3 mol of liquid at same temperature?
a13.0 kcal
b-13.0 kcal
c27.0 kcal
-27.0kcal

Daksh Khanna
8 Years agoGrade 12

1 Answer

Arun
8 Years ago

3H2O(l)→3H2O(g)3H2O(l)→3H2O(g)

Δng=3Δng=3

ΔE=ΔH−ΔngRTΔE=ΔH−ΔngRT

=3×10−3×500×0.002=27Kcal=3×10−3×500×0.002=27Kcal